OpenStudy (anonymous):

Write an equation in slope-intercept form of the line through point P(6, 7) with slope –2.

OpenStudy (anonymous):

y = mx + b plug in slope and the known x, y coordinates to solve for b: 7 = -2(6) + b can you take it from here?

OpenStudy (btaylor):

\[y-y_{1}=m(x-x_{1})\] plug in your given values: \[y-7=-2(x-6)\] simplify and solve for y. You should get \[y = -2x+19\]